OpenStudy (anonymous):

Given two linear equations, explain how you can determine if the lines are parallel without graphing them.

OpenStudy (candyme):

If two lines are parallel, their slopes will be the same. The slope is represented by the m in y=mx+b

OpenStudy (candyme):

For example, if you had y=2x+5 and y=2x+9 These would be parallel because they have the same number in front of the x. But if you had y=4x+3 and y=7x+65 They would NOT be parallel because the number in front of the x is different. Keep in mind that to determine if they are parallel your y has to be alone, no numbers in front, so if you had 5y=5x+10 and y=x+8 they WOULD be parallel, because if you divided the first one by 5 to make the y single, you get y=x+2

OpenStudy (anonymous):

A line passes through the point (–3, –4) and its y-intercept is (0, –9). What is the equation of the line that is parallel to the first line and passes through the point (3, –7)?